OpenStudy (anonymous):

Mary buys gum balls at 4 for 14 cents and sells them at 3 for 15 cents. How many gum balls must Mary sell in order to make a profit of $3.00 ? ( please show ur work)

OpenStudy (anonymous):

cost of 1 gum=14/4=3.5 selling price of 1 gum=15/3=5 profit in 1 gum=5-3.5=1.5 cents so mary will sell=(300/1.5)=200 gums to make a profit of $3.00

OpenStudy (anonymous):

Thanks !